Grade 12IIT JEE Entrance Exam

If the sum of first n terms of an A.P.is pn+qn2 then its common difference is ...

Gaurav g w
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
Dear student
 
if n = 1 then S1 = T1 = p + q
 
now S2 = T1 + T2 = 2p + 4q
 
hence T2 = p + 3q
 
hence common difference = T2– T1 = 2Q